Fw: Status RHIC

From: Flemming Videbaek (videbaek@sgs1.hirg.bnl.gov)
Date: Thu Jun 21 2001 - 15:54:27 EDT

  • Next message: Flemming Videbaek: "Fw: Time Meeting 1:30 PM Tuesday Jun 26"

    ------------------------------------------------------
    Flemming Videbaek
    Physics Department
    Brookhaven National Laboratory
    
    tlf: 631-344-4106
    fax 631-344-1334
    e-mail: videbaek@bnl.gov
    ----- Original Message -----
    From: "Yousef Makdisi" <makdisi@bnl.gov>
    To: "Dana Beavis" <Beavis@bnl.gov>; "Christie William" <Christie@bnl.gov>;
    "Frawley Anthony" <frawley@fsuhip.physics.fsu.edu>; "Robert Pak"
    <Robert.Pak@bnl.gov>
    Cc: "Dejan Trbojevic" <trbojevic@bnl.gov>; "Todd Satogata"
    <satogat1@bnl.gov>; "Thomas Roser" <roser@bnl.gov>; "Phil Pile"
    <pile@bnl.gov>; "Derek Lowenstein" <lowenstein@bnl.gov>; "Tom Kirk"
    <tkirk@bnl.gov>; "Tom Ludlam" <Ludlam1@bnl.gov>; "Nick Samios"
    <samios@bnl.gov>; "Satoshi Ozaki" <Ozaki@bnl.gov>; "Yousef Makdisi"
    <makdisi@bnl.gov>
    Sent: Thursday, June 21, 2001 2:24 PM
    Subject: Status
    
    
    > To all,
    >     This is to keep you abreast of the progress on RHIC. Please share it
    > with your constituents as I do not have access to the full distribution
    > lists:
    >
    > Last night, we had two solid shifts during which significant progress
    > was achieved on the Blue ring. Ten out of possible 12 gamma-t power
    > supplies and quads were used to accelerate the beam through transition
    > and to gamma of 70 GeV. The observed beam loss at transition was
    > approximately 20 percent. It is anticipated that further tuning, the use
    > of the radial loop, and the addition of sextupole corrections before and
    > after transition will further reduce the losses. This is planned for
    > tonight. It is also worth noting that we did not experience any
    > quench-link events either.
    >
    > Progress was also made on the DX training as a current of 6200 amps was
    > achieved. Work continues on this front today. The operating current at
    > 100 GeV is 6300 amps. The training goal is 6400 amps. If completed
    > today, acceleration of beam to 100 GeV will be attempted tonight.
    >
    > Work continues on preparing the power supplies of the Yellow ring. The
    > first stage is to provide the capability to circulate beam at injection.
    > If completed today, then we will attempt injection into the Yellow ring
    > tonight.
    >
    > The second stage concerns tuning the supplies for ramping as well as
    > completing the respective gamma-t work. It is expected to continue this
    > tomorrow. Friday day-time access will be available on an
    > as-needed-basis.
    >
    > Regards,
    >         Yousef
    >
    



    This archive was generated by hypermail 2b30 : Thu Jun 21 2001 - 15:49:44 EDT